A baseball slugger hits a pitch and watches the ball fly into the bleachers for a home run, landing h=8.5 m higher than it was struck. When visiting with the fan that caught the ball, he learned the ball was moving with final velocity vf= 39.45 m / s at an angle θf=29.5° below horizontal when caught. Assume the ball encountered no air resistance, and use a Cartesian coordinate system with the origin located at the ball's initial position.

3A Begin Date: 1/22/2019 12:01:00 AM-Due Date: 1/29/2019 11:59:00 PM End (14%) Problem 7: A baseball slugger hits a pitch and watches the ball fly into the bleachers for a home run, landing h . 8.5 m higher than it was struck. when visiting Date: 5/3/2019 5:00:00 PM 1 with the fan that caught the ball, he learned the ball was moving with final velocity v 39.45 m/s at an angle θ-29.5° below horizontal when caught. Assume the ball encountered no air resistance, and use a Cartesian coordinate system with the origin located at the balls initial position. 0, Ctheexpertta.com 25% Part (a) Create an expression Poiential 100% Submissions 4 5 6 Atempts remaining: 6 3% per attempt) sin(φ)一| sin@p 0 Hint I give up! Feedback:0%deduction per feedback Hints: 0% deduction per hint. Hints ム25% Part (b) Calculate the balls initial vertical velocity, voy, in m/s. - 25% Part (c) Calculate the magnitude of the balls initial velocity, vo, in m/s. 25% Part (d) Find the angle in degrees above the horizontal at which the ball left the bat. 2019 Expert TA. LLC
Answer #1

a)

V0x = Vf cos(theta)

b)

V0x = Vf cos(theta)

= 39.45 m/s x cos29.5

= 34.335 m/s

tan(theta) = Vy/Vx

Vy = Vx tan(theta)

Vy = 34.335 m/s  x tan(-29.5)

= -19.426 m/s

From kinematic equation ,Vy^2 =V0y^2 + 2 a S

V0y = sqrt (Vy^2 + 2 a S)

Voy = sqrt ((-19.426)^2 + 2 x 9.80 x 8.5)

=23.32 m/s

Hence, V0y = 23.32 m/s

c)

V0 = Voy/sin(theta)

= (23.32 m/s)/sin29.5

= 47.36 m/s

Hence, V0 = 47.36 m/s

d)

theta = tan- (23.32/34.335) = 34.2 deg

Hence, theta = 34.2 deg